Crisp Cookware's common stock is expected to pay a dividend of $2.5 a share at the end of this year (D1 = $2.50); its beta is 0.95; the risk-free rate is 4.7%; and the market risk premium is 6%. The dividend is expected to grow at some constant rate g, and the stock currently sells for $35 a share. Assuming the market is in equilibrium, what does the market believe will be the stock's price at the end of 3 years (i.e., what is  )? Do not round intermediate steps. Round your answer to the nearest cent.
Answer #1

Solution:

Calculation of cost of equity(ke) as per CAPM

Ke=Risk free rate+beta(Market risk premium

=4.7%+(0.95*6%)

=10.40% or 0.104

Current share price(P0)=$35

D1=$2.50

We know that,

P0=D1/(ke-growth rate)

Growth rate=Ke-(D1/P0)

=0.1040-($2.50/$35)

=0.0326

=3.26%

Thus stock price at the end of 3rd year=P0(1+growth rate)^3

=$35(1+0.0326)^3

=$38.54